OpenStudy (anonymous):

HELP PLZZZZ At the local pet store, zebra fish cost $2.10 each and neon tetras cost $1.85 each. If Mary bought 15 fish for a total cost of $29.75, not including tax, how many of each type of fish did she buy? Fill in the blanks below.

OpenStudy (anonymous):

how many zebra fish and neon tetras

OpenStudy (anonymous):

and ill fan and medal

OpenStudy (btaylor):

You have to make a system. Let Z = number of zebra fish Let N = number of neon tetras. Z+N=15 2.1Z+1.85N=29.75 multiply first one by -1.85: -1.85 Z-1.85 N = -27.75 2.1 Z + 1.85 N = 29.75 ____________________________ .25 Z = 2.00 Z = 8 N = 7

OpenStudy (anonymous):

thxx